Three point charges arranged as shown. q1= 4.0x10-9C, 2 = 5.0x10-9C, q3= -6.0x10-9C. Find the net force on q1 due to q2 and q3.




q2---------- 0.5m ------q1--------- 0.3m -------q3




1
Expert's answer
2022-03-21T12:47:29-0400

Force

"F_1=\frac{k q_1q_2}{r^2}"


"F_1=\frac{9\times10^9\times5\times10^{-9}\times4\times10^{-9}}{0.5^2}=7.2\times10^{-7}N"

Force

"F_2=\frac{k q_1q_3}{r^2}"


"F_2=-\frac{9\times10^9\times6\times10^{-9}\times4\times10^{-9}}{0.3^2}=-24\times10^{-7}N\\|F_2|=24\times10^{-7}N"

Net force

"F_{net}=F_1+F_2"


"F_{net}=(7.2+24)\times10^{-7}N=31.2\times10^{-7}N"